# Pinpointing Important Genetic Variants via A Feature-level Filter and Group Knockoffs

### v2

@@ 5 Discussions
As more genetic variants are sequenced using the rapidly developing whole-genome sequencing technology, it is now possible to explain more diseases’ heritability via large-scale genetic studies. Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we developed a new filter to select important features using group knockoffs. Specifically, we first define a new family of conditional independence hypotheses that allow inference at feature level under group knockoffs construction. Based on theoretical properties of test statistics when group knockoffs are used, we develop the FVG knockoff filter with FDR control. Extensive simulated experiments (with both synthetic data and real-world genetic data) empirically validate the FDR control of our proposed filter. Compared with the existing knockoff filters, our FVG filter is shown to be powerful in identifying important features with higher precision. When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, we identify as many signals as the group knockoff filter with as high precision as the model-$X$ knockoff filter at the same target FDR level $0.10$. Within several variant groups identified by the group knockoff filter, our method precisely distinguishes the variants that contribute more to AD variation.
Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023). For example, in genetic analysis, variant groups with higher resolution are typically smaller and have higher correlations, making it challenging to identify important signals, but easier to interpret if they are identified. Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber and Ramdas,, 2017; Katsevich and Sabatti,, 2019; Sesia et al.,, 2020; Gablenz and Sabatti,, 2023) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure. In addition, it is also of great interest to adopt the proposed filter in causal inference. By elaborately designing group knockoffs that can eliminate confounding effects, the proposed filter can perform causal features selection with provable FDR control when unmeasured confounders exist. We leave these tasks as future works.
Appendix A Proof of Theorem 1
@@
[…]
@@
[…]
@@
Appendix B Proof of Theorem 2
As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candès et al., (2018). In other words, for each $l$, $\text{sign}(W_{j})$’s for those $j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}$ are i.i.d. coin flips conditional on $\{W_{j}|j\in\mathcal{C}_{l}\}$ and $\{\text{\rm sign}(W_{j})|j\in\mathcal{C}_{l}\setminus\mathcal{H}_{0}^{(\text{fg})}\}$. Thus, by Lemma 3 of Katsevich and Sabatti, (2019), we have
$\textbf{E}\left\{\sup_{t>0}\frac{V^{(l)}_{\mathcal{H}_{0},+}(t)}{1+V^{(l)}_{\mathcal{H}_{0},-}(t)}\right\}\leq 1.93,\quad\text{where }\begin{cases}V^{(l)}_{\mathcal{H}_{0},+}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\geq t\},\\ V^{(l)}_{\mathcal{H}_{0},-}(t)=\#\{j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}|W_{j}\leq-t\}.\\ \end{cases}$ (20)
@@ Appendix B Proof of Theorem 2
$=$ $\mathbf{E}\Biggl{\{}\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr{\}}$
$=$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
$\leq$ $\mathbf{E}\Biggl{\{}\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr{\}}$
@@
Appendix C Extensions
C.1 Multiple Knockoffs
Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez and Zou,, 2019). Here we extend the proposed Algorithm 2 to multiple group knockoffs $\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ where original and knockoff features are simultaneously exchangeable at the level of feature groups. That is to say, with the convention that $\textbf{X}=\widetilde{\textbf{X}}^{(0)}$, $\widetilde{\textbf{X}}^{(0)},\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ satisfy
@@ C.1 Multiple Knockoffs
[…]
for any permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\{0,1,\ldots,M\}$. With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (5), we follow He et al., (2021) to compute feature statistics
@@
Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoffs counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez and Zou,, 2019) that
@@ 2. $\diamond$
Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoffs counterpart of Algorithm 2.
@@
Row $B_{1}$ $B_{2}$ $B_{3}$cdots
1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$cdots
[…]
C.2 Filters based on E-values
Another variation relies on the connection between knockoff filter and $e$-values discussed in Ren and Barber, (2024). Following Ren and Barber, (2024)’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4. By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang and Ramdas,, 2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
@@ C.2 Filters based on E-values
Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by (7) of Ren and Barber, (2024), we have
@@
Finally, by Theorem 2 of of Ren and Barber, (2024), as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control. ∎

### v3

@@ 5 Discussions
As more genetic variants are sequenced using the rapidly developing whole-genome sequencing technology, it is now possible to explain more diseases’ heritability via large-scale genetic studies. Noting that existing knockoff methods either lose power or lack informativeness because of strong correlations among variants in Alzheimer’s disease analysis of the EADB-UKBB dataset, we develop a new filter that uses group knockoffs to simultaneously select important groups and perform fine-mapping within each selected group. Specifically, we first define a new family of conditional independence hypotheses that allow inference at feature level under group knockoffs construction. Based on theoretical properties of test statistics when group knockoffs are used, we develop the FVG knockoff filter with FDR control. Utilizing penalized regressions, our approach can efficiently learn the sparse feature importance and refine the catching set within each group by selecting a subset of most promising features. Extensive simulated experiments with real-world genetic data empirically validate the FDR control of our proposed filter. Compared with the existing knockoff filters, our FVG filter is shown to return small and pure catching sets without missing many important features. When applying the proposed method to AD analysis of the real-world EADB–UKBB dataset, our FVG filter achieves the balance between power (identifies as many signals as the most powerful group knockoff filter) and precision of catching sets (with small average size and high average purity).
Simultaneous inference of conditional independence at different layers of groups is of great necessity as we need to balance the power and the resolution (informativeness) of groups (Katsevich2019; Sesia2020; Gablenz2024). For example, in genetic analysis, variant groups with higher resolution are typically smaller and have higher correlations, making it challenging to identify important signals, but easier to interpret if they are identified. Thus, it is interesting to incorporate the proposed filter into the multilayer testing framework (Barber2017; Katsevich2019; Sesia2020; Gablenz2024) to perform simultaneous inference of $H^{(\text{fg})}$’s at multiple layers of group structure. In addition, it is also of great interest to adopt the proposed filter in causal inference. By elaborately designing group knockoffs that can eliminate confounding effects, the proposed filter can perform causal features selection with provable FDR control when unmeasured confounders exist. We leave these tasks as future works.
Appendix A Supplementary Figures for Section 1.2
(a) Manhattan plot of feature importance statistics of different genetic variants under KnockoffZoom
(2 layers) at the feature-versus-feature level.
(b) Manhattan plot of the contribution of different variants to the importance statistics of the group under KnockoffZoom (3 layers) at the intermediate layer.
(c) Manhattan plot of feature importance statistics of different genetic variants under KnockoffZoom
(3 layers) at the feature-versus-feature level.
Figure 7: Manhanttan plot under KnockoffZoom (2 layers) and KnockoffZoom (3 layers) in a high-LD region between positions 31571218 $\sim$ 32682663, chromosome 6. Here we exclude plots at the group-versus-group level that is the same as Figure 1 (c).
Appendix B Proof of Theorem 1
@@
    ${\mathop{=\joinrel=}^{D}}$ $([\widetilde{\mathbb{X}}_{j},\widetilde{\mathbb{X}}_{B_{1}\setminus\{j\}},{\mathbb{X}}_{j},{\mathbb{X}}_{B_{1}\setminus\{j\}},\{\mathbb{X}_{-B_{1}},\widetilde{\mathbb{X}}_{-B_{1}}\}],\textbf{y}).$
@@
    ${\mathop{=\joinrel=}^{D}}$ $({\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{2}},{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
    ${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{2}},\widetilde{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
    ${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}_{B_{1}},{\textbf{X}}_{B_{1}},\widetilde{\textbf{X}}_{B_{2}},{\textbf{X}}_{B_{2}})|(\textbf{X}_{-(B_{1}\cup B_{2})},\widetilde{\textbf{X}}_{-(B_{1}\cup B_{2})})$
@@
Appendix C Proof of Theorem 2
As feature statistics $W_{j}$’s satisfy the between-group coin-flipping property, we have that $W_{j}$’s within each row of Table 2 satisfy the coin-flipping property introduced in Candes2018. In other words, for each $l$, $\text{sign}(W_{j})$’s for those $j\in\mathcal{C}_{l}\cap\mathcal{H}_{0}^{(\text{fg})}$ are i.i.d. coin flips conditional on $\{W_{j}|j\in\mathcal{C}_{l}\}$ and $\{\text{\rm sign}(W_{j})|j\in\mathcal{C}_{l}\setminus\mathcal{H}_{0}^{(\text{fg})}\}$. Thus, by Lemma 3 of Katsevich2019, we have
[…]
@@ Appendix C Proof of Theorem 2
$=$ $\mathbf{E}\Biggl\{\frac{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\Biggr\}$
$=$ $\mathbf{E}\Biggl\{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot\frac{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}{1\vee(\sum_{l}\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\})}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr\}$
$\leq$ $\mathbf{E}\Biggl\{\sum_{l=1}^{\infty}\text{I}\left(\max_{j\in\mathcal{C}_{l}}|W_{j}|\geq t^{(l)}\right)\cdot\frac{1+\#\{j\in\mathcal{C}_{l}|W_{j}\leq-t^{(l)}\}}{1\vee[\sum_{l}\#\{j\in\mathcal{C}_{l}|W_{j}\geq t^{(l)}\}]}\cdot\frac{V^{(l)}_{\mathcal{H}_{0},+}(t^{(l)})}{1+V^{(l)}_{\mathcal{H}_{0},-}(t^{(l)})}\Biggr\}$
@@
Appendix D Comparisons between the Naive Filter and the FVG Filter
In this section, we empirically compare the naive filter (Algorithm 1), the FVG filter (Algorithm 2) with and without the $1.93$ factor using simulated datasets under the setting of Section 3.1. Over 1000 simulated datasets of different sample sizes, the empirical FDR and power under different target FDR levels are reported in Figure 8. Although both the naive filter and the FVG filter without the 1.93 factor does not possess theoretically rigorous FDR control, their empirical FDRs remain under control. In addition, they both have higher power than the FVG filter (Algorithm 2) with the $1.93$ factor.
Refer to captionRefer to caption
(a) $n=500$.
Refer to captionRefer to caption
(b) $n=1000$.
Refer to captionRefer to caption
(c) $n=2000$.
Figure 8: Empirical FDR and power of the naive filter (Algorithm 1), the FVG filter (Algorithm 2) with and without the $1.93$ factor with respect to the target FDR level ($\alpha$) over 1000 simulated genetic datasets of different sample sizes.
Appendix E Extensions
E.1 Multiple Knockoffs
Although Algorithm 2 with group knockoffs can obtain the rejection set with FDR control, randomness in group knockoffs generation could produce greatly different $\mathcal{R}^{(\text{fg})}$ in different runs, especially in the case that the number of false $H_{j}^{(\text{fg})}$’s is small and Algorithm 2 could return an empty rejection set (Gimenez2019b). Here we extend the proposed Algorithm 2 to multiple group knockoffs $\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ where original and knockoff features are simultaneously exchangeable at the level of feature groups. That is to say, with the convention that $\textbf{X}=\widetilde{\textbf{X}}^{(0)}$, $\widetilde{\textbf{X}}^{(0)},\widetilde{\textbf{X}}^{(1)},\ldots,\widetilde{\textbf{X}}^{(M)}$ satisfy
@@ E.1 Multiple Knockoffs
${\mathop{=\joinrel=}^{D}}$ $(\widetilde{\textbf{X}}^{(0)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(0)}_{B_{K}},\widetilde{\textbf{X}}^{(1)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(1)}_{B_{K}},\ldots,\widetilde{\textbf{X}}^{(M)}_{B_{1}},\ldots,\widetilde{\textbf{X}}^{(M)}_{B_{K}})$
for any permutations $\sigma_{1},\ldots,\sigma_{K}$ of $\{0,1,\ldots,M\}$. With importance scores $\{T_{j}^{(m)}|j=1,\ldots,p;m=0,\ldots,M\}$ obtained in an analogous way of (6), we follow He2021 to compute feature statistics
@@
Specifically, $\kappa_{j}$ and $\tau_{j}$ are the multiple group knockoff counterparts of $\text{sign}(W_{j})$ and $|W_{j}|$ respectively (Gimenez2019b) that
@@ 2. $\diamond$
Based on $\kappa_{j}$’s and $\tau_{j}$’s, we provide Algorithm 3 as the multiple knockoff counterpart of Algorithm 2.
@@
Row $B_{1}$ $B_{2}$ $B_{3}$ $\cdots$
1 $(\kappa_{(11)},\tau_{(11)})$ $(\kappa_{(21)},\tau_{(21)})$ $(\kappa_{(31)},\tau_{(31)})$ $\cdots$
2 $(\kappa_{(12)},\tau_{(12)})$ $(\kappa_{(22)},\tau_{(22)})$ $(\kappa_{(32)},\tau_{(32)})$ $\cdots$
3 $(\kappa_{(13)},\tau_{(13)})$ $(\kappa_{(23)},\tau_{(23)})$ $(\kappa_{(33)},\tau_{(33)})$ $\cdots$
4 $(\kappa_{(14)},\tau_{(14)})$ $(\kappa_{(24)},\tau_{(24)})$ $(\kappa_{(34)},\tau_{(34)})$ $\cdots$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\ddots$
E.2 Filter based on E-values
Another variation relies on the connection between the knockoff filter and $e$-values discussed in Ren2024. Following Ren2024’s way of computing $e$-values of a set of hypotheses whose $W_{j}$’s possess the coin-flipping property, we can analogously compute $e$-values of hypotheses whose $W_{j}$’s are aligned in the same row of Table 2 as shown in steps 4-7 of Algorithm 4. By doing so, an e-value $e_{j}$ is calculated for each hypothesis $H_{j}^{(\text{fg})}$ and we obtain the rejection set $\mathcal{R}^{(\text{fg})}$ by applying the e-BH procedure (Wang2022) to $e$-values $e_{1},\ldots,e_{p}$ (as shown in steps 8-10 of Algorithm 4).
@@ E.2 Filter based on E-values
Because $W_{j}$’s that are in the same row of Table 2 possess the coin-flipping property, by equation (7) of Ren2024, we have
@@
Finally, by Theorem 2 of of Ren2024, as $\sum_{j\in\mathcal{H}_{0}^{(\text{fg})}}\textbf{E}\left\{e_{j}\right\}\leq p$, the e-BH procedure in steps 8-10 of Algorithm 4 returns the rejection set $\mathcal{R}^{(\text{fg})}$ with FDR control. ∎
Appendix F Detailed Definition of Catching Sets of Resolution-Adaptive Approaches
Catching sets of resolution-adaptive approaches are formally defined as follows.
  • $\star$
    For KnockoffZoom (2 layers), based on selection sets $\mathcal{R}_{\text{KZ}}^{(0)},\mathcal{R}_{\text{KZ}}^{(2)}$ at layers 0 and 2, we derive catching sets as
    $\mathcal{CC}_{\text{KZ}}=\mathcal{CC}_{\text{KZ}}^{(0)}\cup\mathcal{CC}_{\text{KZ}}^{(2)},$
    where $\mathcal{CC}_{\text{KZ}}^{(0)}=\{\mathcal{S}_{j,\text{KZ}}^{(0)}=\{j\}|j\in\mathcal{R}_{\text{KZ}}^{(0)}\},$
    $\mathcal{CC}_{\text{KZ}}^{(2)}=\{\mathcal{S}_{k,\text{KZ}}^{(2)}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(2)}\text{ and }j\notin\mathcal{R}_{\text{KZ}}^{(0)}\text{ for all }j\in{B}^{(2)}_{k}\}$.
  • $\star$
    For KnockoffZoom (3 layers), based on selection sets $\mathcal{R}_{\text{KZ}}^{(0)},\mathcal{R}_{\text{KZ}}^{(1)},\mathcal{R}_{\text{KZ}}^{(2)}$ at layers 0, 1 and 2, we derive catching sets as
    $\mathcal{CC}_{\text{KZ}}=\mathcal{CC}_{\text{KZ}}^{(0)}\cup\mathcal{CC}_{\text{KZ}}^{(1)}\cup\mathcal{CC}_{\text{KZ}}^{(2)},$
    where $\mathcal{CC}_{\text{KZ}}^{(0)}=\{\mathcal{S}_{j,\text{KZ}}^{(0)}=\{j\}|j\in\mathcal{R}_{\text{KZ}}^{(0)}\},$
    $\mathcal{CC}_{\text{KZ}}^{(1)}=\{\mathcal{S}_{k,\text{KZ}}^{(1)}={B}^{(1)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(1)}\text{ and }j\notin\mathcal{R}_{\text{KZ}}^{(0)}\text{ for all }j\in{B}^{(1)}_{k}\}$,
    $\mathcal{CC}_{\text{KZ}}^{(2)}=\{\mathcal{S}_{k,\text{KZ}}^{(2)}={B}^{(2)}_{k}|k\in\mathcal{R}_{\text{KZ}}^{(2)}\text{ and }k^{\prime}\notin\mathcal{R}_{\text{KZ}}^{(1)}\text{ for all }{B}^{(1)}_{k^{\prime}}\subset{B}^{(1)}_{k}\}$,
  • $\star$
    For KeLP (2 layers), based on selection sets $\mathcal{R}_{\text{KeLP}}$, we derive catching sets as
    $\mathcal{CC}_{\text{KeLP}}=\mathcal{CC}_{\text{KeLP}}^{(0)}\cup\mathcal{CC}_{\text{KeLP}}^{(2)},$
    where $\mathcal{CC}_{\text{KeLP}}^{(0)}=\{\mathcal{S}_{j,\text{KeLP}}^{(0)}=\{j\}|\{j\}\in\mathcal{R}_{\text{KeLP}}\},$
    $\mathcal{CC}_{\text{KeLP}}^{(2)}=\{\mathcal{S}_{k,\text{KeLP}}^{(2)}={B}^{(2)}_{k}|{B}^{(2)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
  • $\star$
    For KeLP (3 layers), based on selection sets $\mathcal{R}_{\text{KeLP}}$, we derive catching sets as
    $\mathcal{CC}_{\text{KeLP}}=\mathcal{CC}_{\text{KeLP}}^{(0)}\cup\mathcal{CC}_{\text{KeLP}}^{(1)}\cup\mathcal{CC}_{\text{KeLP}}^{(2)},$
    where $\mathcal{CC}_{\text{KeLP}}^{(0)}=\{\mathcal{S}_{j,\text{KeLP}}^{(0)}=\{j\}|\{j\}\in\mathcal{R}_{\text{KeLP}}\},$
    $\mathcal{CC}_{\text{KeLP}}^{(1)}=\{\mathcal{S}_{k,\text{KeLP}}^{(1)}={B}^{(1)}_{k}|{B}^{(1)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
    $\mathcal{CC}_{\text{KeLP}}^{(2)}=\{\mathcal{S}_{k,\text{KeLP}}^{(2)}={B}^{(2)}_{k}|{B}^{(2)}_{k}\in\mathcal{R}_{\text{KeLP}}\}.$
Appendix G Our Implementation of KeLP (Gablenz2024)
We describe our detailed implementation of KeLP (2 layers) under the simulated experiments in Section 3, while the implementation of KeLP (3 layers) is analogous.
As a procedure based on $e$-value, KeLP first compute $e$-values for all $H_{j}^{\text{(ff)}}$’s at the feature-versus-feature level and all $H_{k}^{\text{(gg)}}$’s at the group-versus-group level as follows.
  • For $j=1,\ldots,p$ where $p=1157$, the $e$-value $e_{j}^{\text{(ff)}}$ of $H_{j}^{\text{(ff)}}$ is computed as
    $e_{j}^{\text{(ff)}}=\frac{c^{(\text{ff})}\times I(W_{j}^{\text{(ff)}}\geq t^{\text{(ff)}}_{\gamma^{\text{(ff)}}})}{1+\#\{j|W_{j}^{\text{(ff)}}\leq-t^{\text{(ff)}}_{\gamma^{\text{(ff)}}}\}},$
    where $W_{j}^{\text{(ff)}}$ is the feature importance score of $H_{j}^{\text{(ff)}}$ computed using model-X knockoff and
    $t^{\text{(ff)}}_{\gamma^{\text{(ff)}}}=\min\Biggl\{t>0\Bigg|\frac{1+\#\{j|W_{j}^{\text{(ff)}}\leq-t\}}{1\vee\#\{j|W_{j}^{\text{(ff)}}\geq t\}}\leq\gamma^{\text{(ff)}}\Biggr\}.$
    In other words, $e_{j}^{\text{(ff)}}$’s are $e$-values computed by model-X knockoff filter under the target FDR level $\gamma^{\text{(ff)}}$.
  • The $e$-value $e_{k}^{\text{(gg)}}$ of $H_{k}^{\text{(gg)}}$ with respect to ${B}^{(2)}_{k}$ is computed as
    $e_{k}^{\text{(gg)}}=\frac{c^{(\text{gg})}\times I(W_{k}^{\text{(gg)}}\geq t^{\text{(gg)}}_{\gamma^{\text{(gg)}}})}{1+\#\{j|W_{k}^{\text{(gg)}}\leq-t^{\text{(gg)}}_{\gamma^{\text{(gg)}}}\}},$
    where $W_{k}^{\text{(gg)}}$ is the feature importance score of $H_{k}^{\text{(gg)}}$ computed using group knockoff with respect to ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ and
    $t^{\text{(gg)}}_{\gamma^{\text{(gg)}}}=\min\Biggl\{t>0\Bigg|\frac{1+\#\{j|W_{k}^{\text{(gg)}}\leq-t\}}{1\vee\#\{j|W_{k}^{\text{(gg)}}\geq t\}}\leq\gamma^{\text{(gg)}}\Biggr\}.$
    In other words, $e_{k}^{\text{(gg)}}$’s are $e$-values computed by group knockoff filter with respect to ${B}^{(2)}_{1},\ldots,{B}^{(2)}_{345}$ under the target FDR level $\gamma^{\text{(gg)}}$.
Based on $e_{j}^{\text{(ff)}}$’s and $e_{k}^{\text{(gg)}}$’s, KeLP returns the resolution-adaptive selection set under the FDR level $\alpha$ via solving the constrained optimization problem
$\max\left(\sum_{j=1}^{1157}x_{j}^{(\text{ff})}+\sum_{k=1}^{345}\frac{x_{k}^{(\text{gg})}}{|{B}^{(2)}_{k}|}\right)$, (22)
s.t. $x_{1}^{(\text{ff})},\ldots,x_{1157}^{(\text{ff})},x_{1}^{(\text{gg})},\ldots,x_{345}^{(\text{gg})}\in\{0,1\},$
[…]
Here, $D_{\text{total}}=1157+345$ equals the total number of $e$-values computed at both levels. Finally, catching sets can be obtained as $\{j|\hat{x}_{j}^{(\text{ff})}=1\}\cup\{{B}^{(2)}_{k}|\hat{x}_{k}^{(\text{gg})}=1\}$ based on the solution $(\hat{x}_{1}^{(\text{ff})},\ldots,\hat{x}_{1157}^{(\text{ff})},\hat{x}_{1}^{(\text{gg})},\ldots,\hat{x}_{345}^{(\text{gg})})$ of (22).
In Gablenz2024, it is said that choice of $c^{(\text{ff})}$, $c^{(\text{gg})}$ and $\gamma^{(\text{ff})}$, $\gamma^{(\text{gg})}$ are left to users. Thus, in our simulation (Section 3), we use $c^{(\text{ff})}=1157$ and $c^{(\text{gg})}=345$ to equal the number of hypotheses in both levels. Suppose $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ is the selection set computed by model-X knockoff filter under the target FDR level $\gamma^{\text{(ff)}}$ and $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$ is the selection set computed by the group knockoff filter under the target FDR level $\gamma^{\text{(gg)}}$, we have that
  • $e_{j}^{\text{(ff)}}\geq\frac{1157}{\gamma^{\text{(ff)}}\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|}$ for all $j\in\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$;
  • $e_{k}^{\text{(gg)}}\geq\frac{345}{\gamma^{\text{(gg)}}\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|}$ for all $k\in\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$.
To avoid the case that no catching set is obtained with nonempty $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ and $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$, we elaborately tune $\gamma^{(\text{ff})}$, $\gamma^{(\text{gg})}$ as follows. Specifically, our target is to include two trivial points in the feasible region of (22), including
  • the point where $x_{j}^{(\text{ff})}=\text{I}(j\in\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}})$ for $j=1,\ldots,1157$ and $x_{k}^{(\text{gg})}=0$ for $k=1,\ldots,345$;
  • the point where $x_{j}^{(\text{ff})}=0$ for $j=1,\ldots,1157$ and $x_{k}^{(\text{gg})}=\text{I}(k\in\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}})$ for $k=1,\ldots,345$.
By doing so, we have nonempty catching sets as long as either $\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}$ or $\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}$ is nonempty.
To achieve this, we require that
$\frac{1157}{\gamma^{\text{(ff)}}\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|}$ $\geq\frac{D_{\text{total}}}{\alpha\times|\mathcal{R}^{\text{(ff)}}_{\gamma^{\text{(ff)}}}|},$
$\frac{345}{\gamma^{\text{(gg)}}\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|}$ $\geq\frac{D_{\text{total}}}{\alpha\times|\mathcal{R}^{\text{(gg)}}_{\gamma^{\text{(gg)}}}|},$
from the third and forth line of (22). This leads to our choice that $\gamma^{\text{(ff)}}=1157/D_{\text{total}}\times\alpha$ and $\gamma^{\text{(gg)}}=345/D_{\text{total}}\times\alpha$. When more than $2$ layers are included in the inference, we analogously implement KeLP with $\gamma$ of each layers as $\alpha$ times ratio between the number of hypotheses in this layer and the total number of hypotheses in all layers and the constant $c$ of each layer as the number of hypotheses in this layer.